# Can Diffusion Models Provide Rigorous Uncertainty Quantification for Bayesian Inverse Problems?

### v1

@@ 4.3 EVALUATION METHODS
Mean Variance CMD MMD Mean Variance CMD MMD
Lang-TU $10.594$ $(1.918,17.898)$ $44.17$ $(33.88,50.87)$ $3.14$ $(1.93,4.27)$ $0.389$ $(0.188,0.589)$ $10.076$ $(1.613,17.952)$ $43.99$ $(33.20,51.19)$ $3.05$ $(1.93,4.26)$ $0.380$ $(0.145,0.585)$
Lang-TC $10.597$ $(2.358,18.025)$ $44.84$ $(35.84,51.02)$ $3.13$ $(2.06,4.28)$ $0.399$ $(0.241,0.593)$ – – – – – – – –
Lang-ODE $10.622$ $(2.201,17.217)$ $46.76$ $(36.76,53.64)$ $3.19$ $(1.98,4.24)$ $0.371$ $(0.188,0.506)$ $10.071$ $(1.465,17.538)$ $45.76$ $(34.98,53.04)$ $3.09$ $(1.98,4.26)$ $0.362$ $(0.157,0.514)$
MAP-TU $\mathbf{0.246}$ $(\mathbf{0.059},\mathbf{0.492})$ $1.29$ $(0.55,2.05)$ $\mathbf{0.08}$ $(\mathbf{0.04},\mathbf{0.13})$ $0.130$ $(0.088,0.168)$ $\mathbf{0.262}$ $(\mathbf{0.073},\mathbf{0.569})$ $1.41$ $(0.61,2.21)$ $\mathbf{0.08}$ $(\mathbf{0.04},\mathbf{0.13})$ $0.128$ $(0.085,0.166)$
MAP-TC $0.261$ $(0.054,0.664)$ $1.51$ $(0.52,2.61)$ $0.09$ $(0.04,0.17)$ $0.133$ $(0.091,0.164)$ – – – – – – – –
MAP-ODE $0.502$ $(0.307,0.719)$ $\mathbf{0.80}$ $(\mathbf{0.34},\mathbf{1.38})$ $0.10$ $(0.06,0.15)$ $0.035$ $(0.022,0.047)$ $0.479$ $(0.324,0.651)$ $\mathbf{0.99}$ $(\mathbf{0.48},\mathbf{1.93})$ $0.10$ $(0.06,0.16)$ $0.035$ $(0.023,0.047)$
RTO-TU $1.156$ $(0.270,2.140)$ $5.04$ $(1.20,8.42)$ $0.34$ $(0.12,0.52)$ $0.016$ $(0.002,0.032)$ $1.017$ $(0.198,2.087)$ $4.05$ $(0.60,7.28)$ $0.29$ $(0.09,0.46)$ $\mathbf{0.013}$ $(\mathbf{0.002},\mathbf{0.031})$
RTO-TC $0.857$ $(0.222,1.922)$ $4.72$ $(0.86,12.46)$ $0.30$ $(0.09,0.56)$ $\mathbf{0.015}$ $(\mathbf{0.001},\mathbf{0.043})$ – – – – – – – –
RTO-ODE $1.200$ $(0.490,2.060)$ $6.24$ $(2.57,9.53)$ $0.38$ $(0.19,0.56)$ $0.074$ $(0.046,0.107)$ $1.061$ $(0.355,2.230)$ $5.10$ $(1.91,7.99)$ $0.32$ $(0.14,0.50)$ $0.067$ $(0.037,0.096)$
Reference $0.068$ $(0.036,0.114)$ $0.22$ $(0.07,0.39)$ $0.02$ $(0.01,0.03)$ $0.001$ $(0.001,0.001)$ $0.068$ $(0.036,0.114)$ $0.22$ $(0.07,0.39)$ $0.02$ $(0.01,0.03)$ $0.001$ $(0.001,0.001)$
Table 2: Stylized inpainting study, high noise regime: Average errors in estimation of the mean and pointwise variance of the posterior distribution, and average central moment discrepancy (CMD) and maximum mean discrepancy (MMD) errors, for different BIPSDA methods. Interdecile ranges are also displayed in parentheses. A reference that shows the average errors between two sets of i.i.d. samples from the ground truth posterior is also included. As can be seen, the Langevin dynamics based approaches lead to poor estimates of the mean and variance of the posterior, while the MAP and RTO based approaches provide relatively strong performance.
@@ 5 RESULTS
Figure 2: Stylized inpainting study, high noise regime: Scatter plots of samples from three different BISPDA methods, as well as ground truth posterior samples, in a representative trial. Two components of the ten-dimensional samples are visualized, with one of the two components in the null space of the forward operator. The samples are overlaid over the posterior density. As can be seen, the RTO and MAP estimation based approaches both provide competitive performance, whereas the Langevin dynamics based approach significantly overestimates the posterior variance.
@@ 5.1 STYLIZED INPAINTING STUDIES
We first examine the performance of the framework in the stylized inpainting studies. Table 1 shows the average error in the low noise regime with regards to the mean and pointwise variance of the posterior, as well as CMD and MMD metrics, for the nine BIPSDA algorithms tested. The corresponding results in the high noise regime are given in Table 2. Here we first note that when the analytic scores are used, the Tweedie-Correlated (TC) based approaches perform comparably with the Tweedie-Uncorrelated (TU) based approaches. However, TC-based approaches were not implemented when using the learned scores, as the Jacobian of the learned score is not a reliable estimator of the denoising distribution covariance matrix. The performance of the TC-based methods in the learned score regime was thus omitted from this and all subsequent studies.
All of the methods that were tested perform fairly well in the low noise regime using the learned score, with the RTO-TU approach providing particularly strong performance. Interestingly, the Lang-ODE approach outperforms Lang-TU despite using a biased estimate of the mean of the denoising distribution (see Appendix 7). In contrast, in the high noise regime, the Langevin dynamics-based approaches all perform significantly worse than the other approaches, while the MAP estimation-based approaches perform well despite lacking the theoretical basis of the Langevin dynamics and RTO-based approaches. Further, this discrepancy is not due to score modeling error, as the Langevin dynamics-based approaches perform poorly even when the analytic scores are used.
Figure 2, which displays scatter plots of samples from three different BIPSDA methods in a representative high-noise trial, provides insight into the failure of the Langevin-based approaches. In particular, it can be seen that the Langevin dynamics based approach produces samples in low-density regions with respect to the likelihood and therefore overestimates the variance of the posterior distribution. The RTO-based approach, while still overestimating the posterior variance, is much more accurate. Finally, the MAP estimation-based approach, while underestimating the variance of each of the modes, provides the most accurate estimate of the weights of the modes of all the methods tested.
@@ 5.2 STYLIZED X-RAY TOMOGRAPHY STUDY
We now analyze the results from the stylized x-ray tomography study. Table 3 reports the average mean, variance, CMD, and MMD errors for the different BIPSDA algorithms in this problem setting. As can be seen, in this problem setting all of the approaches we considered performed fairly well, despite the non-linearity of the forward model. The RTO-TU approach provides particularly strong performance and is the top performing approach with respect to all metrics we considered in the learned score regime. This demonstrates the potential of the RTO based approaches in the context of non-linear inverse problems.
Figure 3 shows scatter plots of samples from the ground truth posterior and the Lang-TU, MAP-TU, and RTO-TU algorithms for a representative trial. As can be seen, for this problem the posterior is approximately unimodal. Further, while all of the tested algorithms perform fairly well, the MAP-TU approach underestimates the posterior variance, which is consistent with the results of the stylized inpainting study. The RTO-TU approach is able to correct the variance underestimation of the MAP-TU approach and performs similarly to the Lang-TU method on this trial.
@@
Mean Variance CMD MMD Mean Variance CMD MMD
Lang-TU $0.44$ $(0.27,0.64)$ $0.22$ $(0.04,0.58)$ $0.12$ $(0.07,0.18)$ $0.016$ $(0.008,0.025)$ $0.44$ $(0.26,0.68)$ $0.25$ $(0.05,0.61)$ $0.12$ $(0.07,0.19)$ $0.017$ $(0.009,0.026)$
Lang-TC $0.45$ $(0.28,0.65)$ $0.23$ $(0.04,0.60)$ $0.12$ $(0.08,0.18)$ $0.017$ $(0.008,0.028)$ – – – – – – – –
Lang-ODE $0.26$ $(0.10,0.48)$ $1.19$ $(0.82,1.57)$ $0.15$ $(0.08,0.23)$ $0.031$ $(0.026,0.036)$ $0.27$ $(0.10,0.51)$ $1.14$ $(0.82,1.43)$ $0.15$ $(0.08,0.22)$ $0.031$ $(0.024,0.037)$
MAP-TU $0.12$ $(0.06,0.25)$ $1.66$ $(1.07,2.20)$ $0.14$ $(0.08,0.21)$ $0.162$ $(0.151,0.174)$ $\mathbf{0.13}$ $(\mathbf{0.06},\mathbf{0.25})$ $1.67$ $(1.06,2.24)$ $0.14$ $(0.08,0.21)$ $0.163$ $(0.149,0.175)$
MAP-TC $\mathbf{0.09}$ $(\mathbf{0.03},\mathbf{0.26})$ $1.65$ $(1.06,2.20)$ $0.13$ $(0.07,0.22)$ $0.160$ $(0.149,0.171)$ – – – – – – – –
MAP-ODE $0.55$ $(0.35,0.75)$ $0.98$ $(0.67,1.26)$ $0.19$ $(0.14,0.25)$ $0.075$ $(0.043,0.112)$ $0.54$ $(0.35,0.70)$ $1.01$ $(0.67,1.33)$ $0.19$ $(0.15,0.25)$ $0.076$ $(0.045,0.113)$
RTO-TU $0.12$ $(0.05,0.27)$ $0.19$ $(0.04,0.58)$ $\mathbf{0.04}$ $(\mathbf{0.02},\mathbf{0.11})$ $0.003$ $(0.001,0.007)$ $\mathbf{0.13}$ $(\mathbf{0.05},\mathbf{0.26})$ $\mathbf{0.19}$ $(\mathbf{0.04},\mathbf{0.55})$ $\mathbf{0.04}$ $(\mathbf{0.02},\mathbf{0.10})$ $\mathbf{0.003}$ $(\mathbf{0.001},\mathbf{0.006})$
RTO-TC $\mathbf{0.09}$ $(\mathbf{0.03},\mathbf{0.27})$ $\mathbf{0.18}$ $(\mathbf{0.04},\mathbf{0.58})$ $\mathbf{0.04}$ $(\mathbf{0.01},\mathbf{0.11})$ $\mathbf{0.002}$ $(\mathbf{0.001},\mathbf{0.007})$ – – – – – – – –
RTO-ODE $0.55$ $(0.36,0.73)$ $1.51$ $(0.85,2.25)$ $0.24$ $(0.16,0.33)$ $0.054$ $(0.043,0.068)$ $0.52$ $(0.34,0.69)$ $1.41$ $(0.84,2.00)$ $0.23$ $(0.16,0.30)$ $0.051$ $(0.038,0.069)$
Reference $0.05$ $(0.04,0.07)$ $0.07$ $(0.04,0.11)$ $0.02$ $(0.01,0.03)$ $0.001$ $(0.001,0.001)$ $0.05$ $(0.04,0.07)$ $0.07$ $(0.04,0.11)$ $0.02$ $(0.01,0.03)$ $0.001$ $(0.001,0.001)$
@@
Table 3: Stylized x-ray tomography study: Average errors in estimation of the mean and pointwise variance of the posterior distribution, and average central moment discrepancy (CMD) and maximum mean discrepancy (MMD) errors, for different BIPSDA methods. Interdecile ranges are also displayed in parentheses. A reference that shows the average errors between two sets of i.i.d. samples from the ground truth posterior is also included. While all methods provide competitive performance, the RTO-TU approach outperforms all other approaches in almost every metric.
@@
Figure 3: Stylized x-ray tomography study: Scatter plots of samples from three different BISPDA methods, as well as ground truth posterior samples, in a representative trial. Here the samples have been projected onto the space spanned by the two right singular vectors corresponding to the largest and smallest singular values of the forward operator $\mathbf{C}$. The samples are overlaid over a kernel density approximation of the pushforward of the posterior density under this projection. As can be seen, all three methods perform well in this setting, with the Langevin dynamics and RTO based approaches providing particularly strong performance.
@@ 5.3 STYLIZED PHASE RETRIEVAL STUDY
Mean Variance CMD MMD Mean Variance CMD MMD
Lang-TU $2.42$ $(0.08,5.88)$ $6.26$ $(1.62,14.25)$ $0.20$ $(0.01,0.46)$ $0.076$ $(0.002,0.152)$ $2.45$ $(0.07,5.93)$ $5.73$ $(0.67,14.61)$ $0.20$ $(0.01,0.46)$ $0.073$ $(0.001,0.157)$
Lang-TC $\mathbf{1.90}$ $(\mathbf{0.05},\mathbf{4.34})$ $\mathbf{4.01}$ $(\mathbf{0.13},\mathbf{9.69})$ $\mathbf{0.15}$ $(\mathbf{0.00},\mathbf{0.34})$ $0.048$ $(0.001,0.064)$ – – – – – – – –
Lang-ODE $2.53$ $(0.11,5.95)$ $8.33$ $(3.60,15.93)$ $0.21$ $(0.02,0.47)$ $0.134$ $(0.049,0.230)$ $2.54$ $(0.08,5.92)$ $7.85$ $(2.47,16.20)$ $0.21$ $(0.01,0.47)$ $0.128$ $(0.044,0.217)$
MAP-TU $1.97$ $(0.06,4.37)$ $5.88$ $(1.41,11.63)$ $0.16$ $(0.01,0.35)$ $0.129$ $(0.064,0.177)$ $\mathbf{1.97}$ $(\mathbf{0.08},\mathbf{4.42})$ $\mathbf{5.63}$ $(\mathbf{1.32},\mathbf{11.46})$ $\mathbf{0.16}$ $(\mathbf{0.01},\mathbf{0.36})$ $0.127$ $(0.064,0.169)$
MAP-TC $1.99$ $(0.04,4.74)$ $4.82$ $(1.41,10.29)$ $0.16$ $(0.01,0.38)$ $0.129$ $(0.072,0.150)$ – – – – – – – –
MAP-ODE $2.05$ $(0.07,4.72)$ $6.55$ $(2.09,13.57)$ $0.17$ $(0.02,0.38)$ $0.069$ $(0.015,0.101)$ $2.06$ $(0.08,4.75)$ $6.24$ $(1.54,12.68)$ $0.17$ $(0.01,0.38)$ $0.070$ $(0.015,0.103)$
RTO-TU $2.53$ $(0.11,6.12)$ $6.83$ $(2.44,13.11)$ $0.21$ $(0.02,0.47)$ $0.069$ $(0.006,0.165)$ $2.53$ $(0.09,6.11)$ $5.98$ $(1.59,12.47)$ $0.20$ $(0.01,0.47)$ $\mathbf{0.065}$ $(\mathbf{0.003},\mathbf{0.156})$
RTO-TC $2.04$ $(0.05,4.85)$ $4.68$ $(0.20,11.29)$ $0.16$ $(0.00,0.37)$ $\mathbf{0.047}$ $(\mathbf{0.001},\mathbf{0.090})$ – – – – – – – –
RTO-ODE $2.57$ $(0.16,5.94)$ $8.64$ $(4.03,14.63)$ $0.22$ $(0.03,0.47)$ $0.124$ $(0.055,0.210)$ $2.58$ $(0.12,5.99)$ $7.71$ $(3.15,13.69)$ $0.21$ $(0.02,0.47)$ $0.117$ $(0.048,0.207)$
Reference $0.24$ $(0.05,0.50)$ $0.79$ $(0.08,1.15)$ $0.02$ $(0.00,0.04)$ $0.002$ $(0.001,0.002)$ $0.24$ $(0.05,0.50)$ $0.79$ $(0.08,1.15)$ $0.02$ $(0.00,0.04)$ $0.002$ $(0.001,0.002)$
@@
Table 4: Stylized phase retrieval study: Average errors in estimation of the mean and pointwise variance of the posterior distribution, and average central moment discrepancy (CMD) and maximum mean discrepancy (MMD) errors, for different BIPSDA methods. Interdecile ranges are also displayed in parentheses, and a reference that shows the average errors between two sets of i.i.d. samples from the ground truth posterior is included. Note that for the Lang-TU and RTO-TU methods in the learned score regime, 29 and 185 samples, respectively, of the million total samples generated across the hundred trials numerically diverged and were discarded. As can be seen, all of the algorithms we analyzed produce significant errors in posterior sampling on this challenging problem.
@@
Figure 4: Stylized phase retrieval study: Scatter plots of samples from three different BISPDA methods, as well as ground truth posterior samples, in a representative trial. Here the samples have been projected onto onto the space spanned by the two right singular vectors corresponding to the largest and smallest singular values of the forward operator $\mathbf{B}$. The samples are overlaid over a kernel density approximation of the pushforward of the posterior density under this projection. As can be seen, the RTO and Langevin based approaches both abandon samples between the two distribution modes, while the MAP approach underestimates the variance of the modes.
@@
We now examine the performance of the algorithms in the stylized phase retrieval study. Figure 4 shows scatter plots of samples from the ground-truth posterior and the Lang-TU, MAP-TU, and RTO-TU algorithms in a representative trial (with learned score). As can be seen, both the Lang-TU and RTO-TU algorithms produce samples that lie in a low-density region between the two posterior modes. The MAP-TU approach avoids this issue, but suffers from significant underestimation of the variance of each of the modes.

### v2

@@ 4.3 EVALUATION METHODS
Mean Error Variance Error CMD MMD Mean Error Variance Error CMD MMD
Lang-TU
$10.594$
$(1.918,17.898)$
$44.17$
$(33.88,50.87)$
$3.14$
$(1.93,4.27)$
$0.389$
$(0.188,0.589)$
$10.076$
$(1.613,17.952)$
$43.99$
$(33.20,51.19)$
$3.05$
$(1.93,4.26)$
$0.380$
$(0.145,0.585)$
Lang-TC
$10.597$
$(2.358,18.025)$
$44.84$
$(35.84,51.02)$
$3.13$
$(2.06,4.28)$
$0.399$
$(0.241,0.593)$
Lang-ODE
$10.622$
$(2.201,17.217)$
$46.76$
$(36.76,53.64)$
$3.19$
$(1.98,4.24)$
$0.371$
$(0.188,0.506)$
$10.071$
$(1.465,17.538)$
$45.76$
$(34.98,53.04)$
$3.09$
$(1.98,4.26)$
$0.362$
$(0.157,0.514)$
MAP-TU
$\mathbf{0.246}$
$(\mathbf{0.059},\mathbf{0.492})$
$1.29$
$(0.55,2.05)$
$\mathbf{0.08}$
$(\mathbf{0.04},\mathbf{0.13})$
$0.130$
$(0.088,0.168)$
$\mathbf{0.262}$
$(\mathbf{0.073},\mathbf{0.569})$
$1.41$
$(0.61,2.21)$
$\mathbf{0.08}$
$(\mathbf{0.04},\mathbf{0.13})$
$0.128$
$(0.085,0.166)$
MAP-TC
$0.261$
$(0.054,0.664)$
$1.51$
$(0.52,2.61)$
$0.09$
$(0.04,0.17)$
$0.133$
$(0.091,0.164)$
MAP-ODE
$0.502$
$(0.307,0.719)$
$\mathbf{0.80}$
$(\mathbf{0.34},\mathbf{1.38})$
$0.10$
$(0.06,0.15)$
$0.035$
$(0.022,0.047)$
$0.479$
$(0.324,0.651)$
$\mathbf{0.99}$
$(\mathbf{0.48},\mathbf{1.93})$
$0.10$
$(0.06,0.16)$
$0.035$
$(0.023,0.047)$
RTO-TU
$1.156$
$(0.270,2.140)$
$5.04$
$(1.20,8.42)$
$0.34$
$(0.12,0.52)$
$0.016$
$(0.002,0.032)$
$1.017$
$(0.198,2.087)$
$4.05$
$(0.60,7.28)$
$0.29$
$(0.09,0.46)$
$\mathbf{0.013}$
$(\mathbf{0.002},\mathbf{0.031})$
RTO-TC
$0.857$
$(0.222,1.922)$
$4.72$
$(0.86,12.46)$
$0.30$
$(0.09,0.56)$
$\mathbf{0.015}$
$(\mathbf{0.001},\mathbf{0.043})$
RTO-ODE
$1.200$
$(0.490,2.060)$
$6.24$
$(2.57,9.53)$
$0.38$
$(0.19,0.56)$
$0.074$
$(0.046,0.107)$
$1.061$
$(0.355,2.230)$
$5.10$
$(1.91,7.99)$
$0.32$
$(0.14,0.50)$
$0.067$
$(0.037,0.096)$
Reference
$0.068$
$(0.036,0.114)$
$0.22$
$(0.07,0.39)$
$0.02$
$(0.01,0.03)$
$0.001$
$(0.001,0.001)$
$0.068$
$(0.036,0.114)$
$0.22$
$(0.07,0.39)$
$0.02$
$(0.01,0.03)$
$0.001$
$(0.001,0.001)$
Table 3: Stylized inpainting study, high noise regime: Average errors in estimation of the mean and pointwise variance of the posterior distribution, and average central moment discrepancy (CMD) and maximum mean discrepancy (MMD) errors, for different BIPSDA methods. Interdecile ranges are also displayed in parentheses. A reference that shows the average errors between two sets of i.i.d. samples from the ground truth posterior is also included. As can be seen, the ‘Lang’ variants lead to poor estimates of the mean and variance of the posterior, while the ‘MAP’ and ‘RTO’ variants provide relatively strong performance.
@@ 5 RESULTS
Figure 2: Stylized inpainting study, high noise regime: Scatter plots of samples from three different BIPSDA methods, as well as ground truth posterior samples, in a representative trial. Two components of the ten-dimensional samples are visualized, with one of the two components in the null space of the forward operator. The samples are overlaid over the posterior density. As can be seen, the ‘RTO’ and ‘MAP’ variants both provide competitive performance, whereas the ‘Lang’ variant significantly overestimates the posterior variance.
@@ 5.1 STYLIZED INPAINTING STUDIES
We first examine the performance of the framework in the stylized inpainting studies. Table 2 shows the average error in the low noise regime with regards to the mean and pointwise variance of the posterior, as well as CMD and MMD metrics, for the nine BIPSDA algorithms tested. The corresponding results in the high noise regime are given in Table 3. Here we first note that when the analytic scores are used, the Tweedie-Correlated (‘TC’) based approaches perform comparably with the Tweedie-Uncorrelated (‘TU’) based approaches. However, ‘TC’ variants were not implemented when using the learned scores, as the Jacobian of the learned score is not a reliable estimator of the denoising distribution covariance matrix. The performance of the ‘TC’ variants in the learned score regime was thus omitted from this and all subsequent studies.
All of the methods that were tested perform fairly well in the low noise regime using the learned score, with the ‘RTO-TU’ approach providing particularly strong performance. In contrast, in the high noise regime, the Langevin dynamics-based approaches all perform significantly worse than the other approaches, while the ‘MAP’ variants perform well despite lacking the theoretical basis of the ‘Lang’ and ‘RTO’ variants. Further, this discrepancy is not due to score modeling error, as the ‘Lang’ variants perform poorly even when the analytic scores are used.
Figure 2, which displays scatter plots of samples from three different BIPSDA methods in a representative high-noise trial, provides insight into the failure of the ‘Lang’ variants. In particular, it can be seen that the Langevin dynamics based approach produces samples in low-density regions with respect to the likelihood and therefore overestimates the variance of the posterior distribution. The ‘RTO’ variants, while still overestimating the posterior variance, are much more accurate. Finally, the ‘MAP’ variants, while underestimating the variance of each of the modes, provide the most accurate estimate of the weights of the modes of all the methods tested.
@@ 5.2 STYLIZED X-RAY TOMOGRAPHY STUDY
We now analyze the results from the stylized x-ray tomography study. Table 4 reports the average mean, variance, CMD, and MMD errors for the different BIPSDA algorithms in this problem setting. As can be seen, in this problem setting all of the approaches we considered performed fairly well, despite the non-linearity of the forward model. The ‘RTO-TU’ approach provides particularly strong performance and is the top performing approach with respect to all metrics we considered in the learned score regime. This demonstrates the potential of the RTO-based approaches in the context of non-linear inverse problems.
Figure 3 shows scatter plots of samples from the ground truth posterior and the ‘Lang-TU’, ‘MAP-TU’, and ‘RTO-TU’ algorithms for a representative trial. As can be seen, for this problem the posterior is approximately unimodal. Further, while all of the tested algorithms perform fairly well, the ‘MAP-TU’ approach underestimates the posterior variance, which is consistent with the results of the stylized inpainting study. The ‘RTO-TU’ approach is able to correct the variance underestimation of the ‘MAP-TU’ approach and performs similarly to the ‘Lang-TU’ method on this trial.
@@
Mean Error Variance Error CMD MMD Mean Error Variance Error CMD MMD
Lang-TU
$0.44$
$(0.27,0.64)$
$0.22$
$(0.04,0.58)$
$0.12$
$(0.07,0.18)$
$0.016$
$(0.008,0.025)$
$0.44$
$(0.26,0.68)$
$0.25$
$(0.05,0.61)$
$0.12$
$(0.07,0.19)$
$0.017$
$(0.009,0.026)$
Lang-TC
$0.45$
$(0.28,0.65)$
$0.23$
$(0.04,0.60)$
$0.12$
$(0.08,0.18)$
$0.017$
$(0.008,0.028)$
Lang-ODE
$0.26$
$(0.10,0.48)$
$1.19$
$(0.82,1.57)$
$0.15$
$(0.08,0.23)$
$0.031$
$(0.026,0.036)$
$0.27$
$(0.10,0.51)$
$1.14$
$(0.82,1.43)$
$0.15$
$(0.08,0.22)$
$0.031$
$(0.024,0.037)$
MAP-TU
$0.12$
$(0.06,0.25)$
$1.66$
$(1.07,2.20)$
$0.14$
$(0.08,0.21)$
$0.162$
$(0.151,0.174)$
$\mathbf{0.13}$
$(\mathbf{0.06},\mathbf{0.25})$
$1.67$
$(1.06,2.24)$
$0.14$
$(0.08,0.21)$
$0.163$
$(0.149,0.175)$
MAP-TC
$\mathbf{0.09}$
$(\mathbf{0.03},\mathbf{0.26})$
$1.65$
$(1.06,2.20)$
$0.13$
$(0.07,0.22)$
$0.160$
$(0.149,0.171)$
MAP-ODE
$0.55$
$(0.35,0.75)$
$0.98$
$(0.67,1.26)$
$0.19$
$(0.14,0.25)$
$0.075$
$(0.043,0.112)$
$0.54$
$(0.35,0.70)$
$1.01$
$(0.67,1.33)$
$0.19$
$(0.15,0.25)$
$0.076$
$(0.045,0.113)$
RTO-TU
$0.12$
$(0.05,0.27)$
$0.19$
$(0.04,0.58)$
$\mathbf{0.04}$
$(\mathbf{0.02},\mathbf{0.11})$
$0.003$
$(0.001,0.007)$
$\mathbf{0.13}$
$(\mathbf{0.05},\mathbf{0.26})$
$\mathbf{0.19}$
$(\mathbf{0.04},\mathbf{0.55})$
$\mathbf{0.04}$
$(\mathbf{0.02},\mathbf{0.10})$
$\mathbf{0.003}$
$(\mathbf{0.001},\mathbf{0.006})$
RTO-TC
$\mathbf{0.09}$
$(\mathbf{0.03},\mathbf{0.27})$
$\mathbf{0.18}$
$(\mathbf{0.04},\mathbf{0.58})$
$\mathbf{0.04}$
$(\mathbf{0.01},\mathbf{0.11})$
$\mathbf{0.002}$
$(\mathbf{0.001},\mathbf{0.007})$
RTO-ODE
$0.55$
$(0.36,0.73)$
$1.51$
$(0.85,2.25)$
$0.24$
$(0.16,0.33)$
$0.054$
$(0.043,0.068)$
$0.52$
$(0.34,0.69)$
$1.41$
$(0.84,2.00)$
$0.23$
$(0.16,0.30)$
$0.051$
$(0.038,0.069)$
Reference
$0.05$
$(0.04,0.07)$
$0.07$
$(0.04,0.11)$
$0.02$
$(0.01,0.03)$
$0.001$
$(0.001,0.001)$
$0.05$
$(0.04,0.07)$
$0.07$
$(0.04,0.11)$
$0.02$
$(0.01,0.03)$
$0.001$
$(0.001,0.001)$
@@
Table 4: Stylized x-ray tomography study: Average errors in estimation of the mean and pointwise variance of the posterior distribution, and average central moment discrepancy (CMD) and maximum mean discrepancy (MMD) errors, for different BIPSDA methods. Interdecile ranges are also displayed in parentheses. A reference that shows the average errors between two sets of i.i.d. samples from the ground truth posterior is also included. While all methods provide competitive performance, the ‘RTO-TU’ approach outperforms all other approaches in almost every metric.
@@
Figure 3: Stylized x-ray tomography study: Scatter plots of samples from three different BIPSDA methods, as well as ground truth posterior samples, in a representative trial. Here the samples have been projected onto the space spanned by the two right singular vectors corresponding to the largest and smallest singular values of the forward operator $\mathbf{C}$. The samples are overlaid over a kernel density approximation of the pushforward of the posterior density under this projection. As can be seen, all three methods perform well in this setting, with the ‘Lang’ and ‘RTO’ variants providing particularly strong performance.
@@ 5.3 STYLIZED PHASE RETRIEVAL STUDY
Mean Error Variance Error CMD MMD Mean Error Variance Error CMD MMD
Lang-TU
$2.42$
$(0.08,5.88)$
$6.26$
$(1.62,14.25)$
$0.20$
$(0.01,0.46)$
$0.076$
$(0.002,0.152)$
$2.45$
$(0.07,5.93)$
$5.73$
$(0.67,14.61)$
$0.20$
$(0.01,0.46)$
$0.073$
$(0.001,0.157)$
Lang-TC
$\mathbf{1.90}$
$(\mathbf{0.05},\mathbf{4.34})$
$\mathbf{4.01}$
$(\mathbf{0.13},\mathbf{9.69})$
$\mathbf{0.15}$
$(\mathbf{0.00},\mathbf{0.34})$
$0.048$
$(0.001,0.064)$
Lang-ODE
$2.53$
$(0.11,5.95)$
$8.33$
$(3.60,15.93)$
$0.21$
$(0.02,0.47)$
$0.134$
$(0.049,0.230)$
$2.54$
$(0.08,5.92)$
$7.85$
$(2.47,16.20)$
$0.21$
$(0.01,0.47)$
$0.128$
$(0.044,0.217)$
MAP-TU
$1.97$
$(0.06,4.37)$
$5.88$
$(1.41,11.63)$
$0.16$
$(0.01,0.35)$
$0.129$
$(0.064,0.177)$
$\mathbf{1.97}$
$(\mathbf{0.08},\mathbf{4.42})$
$\mathbf{5.63}$
$(\mathbf{1.32},\mathbf{11.46})$
$\mathbf{0.16}$
$(\mathbf{0.01},\mathbf{0.36})$
$0.127$
$(0.064,0.169)$
MAP-TC
$1.99$
$(0.04,4.74)$
$4.82$
$(1.41,10.29)$
$0.16$
$(0.01,0.38)$
$0.129$
$(0.072,0.150)$
MAP-ODE
$2.05$
$(0.07,4.72)$
$6.55$
$(2.09,13.57)$
$0.17$
$(0.02,0.38)$
$0.069$
$(0.015,0.101)$
$2.06$
$(0.08,4.75)$
$6.24$
$(1.54,12.68)$
$0.17$
$(0.01,0.38)$
$0.070$
$(0.015,0.103)$
RTO-TU
$2.53$
$(0.11,6.12)$
$6.83$
$(2.44,13.11)$
$0.21$
$(0.02,0.47)$
$0.069$
$(0.006,0.165)$
$2.53$
$(0.09,6.11)$
$5.98$
$(1.59,12.47)$
$0.20$
$(0.01,0.47)$
$\mathbf{0.065}$
$(\mathbf{0.003},\mathbf{0.156})$
RTO-TC
$2.04$
$(0.05,4.85)$
$4.68$
$(0.20,11.29)$
$0.16$
$(0.00,0.37)$
$\mathbf{0.047}$
$(\mathbf{0.001},\mathbf{0.090})$
RTO-ODE
$2.57$
$(0.16,5.94)$
$8.64$
$(4.03,14.63)$
$0.22$
$(0.03,0.47)$
$0.124$
$(0.055,0.210)$
$2.58$
$(0.12,5.99)$
$7.71$
$(3.15,13.69)$
$0.21$
$(0.02,0.47)$
$0.117$
$(0.048,0.207)$
Reference
$0.24$
$(0.05,0.50)$
$0.79$
$(0.08,1.15)$
$0.02$
$(0.00,0.04)$
$0.002$
$(0.001,0.002)$
$0.24$
$(0.05,0.50)$
$0.79$
$(0.08,1.15)$
$0.02$
$(0.00,0.04)$
$0.002$
$(0.001,0.002)$
@@
Table 5: Stylized phase retrieval study: Average errors in estimation of the mean and pointwise variance of the posterior distribution, and average central moment discrepancy (CMD) and maximum mean discrepancy (MMD) errors, for different BIPSDA methods. Interdecile ranges are also displayed in parentheses, and a reference that shows the average errors between two sets of i.i.d. samples from the ground truth posterior is included. Note that for the ‘Lang-TU’ and ‘RTO-TU’ methods in the learned score regime, 29 and 185 samples, respectively, of the million total samples generated across the hundred trials numerically diverged and were discarded. As can be seen, all of the algorithms we analyzed produce significant errors in posterior sampling on this challenging problem.
@@
Figure 4: Stylized phase retrieval study: Scatter plots of samples from three different BIPSDA methods, as well as ground truth posterior samples, in a representative trial. Here the samples have been projected onto onto the space spanned by the two right singular vectors corresponding to the largest and smallest singular values of the forward operator $\mathbf{B}$. The samples are overlaid over a kernel density approximation of the pushforward of the posterior density under this projection. As can be seen, the ‘RTO’ and ‘Lang’ variants both abandon samples between the two distribution modes, while the ‘MAP’ variant underestimates the variance of the modes.
@@
We now examine the performance of the algorithms in the stylized phase retrieval study. Figure 4 shows scatter plots of samples from the ground-truth posterior and the ‘Lang-TU’, ‘MAP-TU’, and ‘RTO-TU’ algorithms in a representative trial (with learned score). As can be seen, both the ‘Lang-TU’ and ‘RTO-TU’ algorithms produce samples that lie in a low-density region between the two posterior modes. The ‘MAP-TU’ approach avoids this issue, but suffers from significant underestimation of the variance of each of the modes.